OpenStudy (anonymous):

90/5+4*2-1=19? how do I do this?!

OpenStudy (anonymous):

take it daniel and rsvitale

OpenStudy (anonymous):

90/5+4*2-1 is equal to 25 not 19. 90/5=18, (4*2)-1=7, 18+7=25. Can you tell us what the question was asking exactly?

OpenStudy (anonymous):

GUYS I NEED AN ANSWER FAST!

OpenStudy (anonymous):

90/(5+4) = 10 10*2 -1 =19

OpenStudy (anonymous):

Thats the question, I have to figure out that equation...

OpenStudy (anonymous):

oh you need to use parenthesis Mia

OpenStudy (anonymous):

Yeah Can you help?

OpenStudy (anonymous):

it should be written ((90/(5+4))*2)-1. This is equal to 19. I used more parenthesis than necessary to be clear.

OpenStudy (anonymous):

Thanks SOOO much

OpenStudy (anonymous):

I need help now on solving it

OpenStudy (anonymous):

ok first we add 5 and 4 to get 9. so 90/(5+4)=90/9=10. Then we take that piece and multiply it by 2. 2*10=20. Finally we subtract 1. 20-1=19

OpenStudy (anonymous):

the parenthesis indicate order of operations

OpenStudy (anonymous):

does that make sense or do you have any questions?

OpenStudy (anonymous):

Yep, that makes perfect sense... um lemme check hang on

OpenStudy (anonymous):

By inserting one pair of parenthesis at different places in 100-3^2+5*4-6/2 find as many different ways of evaluating the expression as you can.

OpenStudy (anonymous):

ok i think there are many ways to get a different answer. (100-3)^2+5*4-6/2, 100-3^2+(5*4-6)/2, (100-3^2+5*4-6)/2, are several groupings that give different answers. try to think of some more.

OpenStudy (anonymous):

the parenthesis make a big difference; for example 100-3^2=91 but (100-3)^2=9409

OpenStudy (anonymous):

Oh yeah....i see..

OpenStudy (anonymous):

how many more ways are there?

OpenStudy (anonymous):

that depends on how it is written. Is the 3^2 actually written with a carat "^" symbol or just as a superscript

OpenStudy (anonymous):

there are a bunch of ways to do it if we are just writing it as it's typed but if it is written a certain way in a text book there are less possibilities I would have to know exactly how it was written

OpenStudy (anonymous):

the exponent is up in the right corner of the "3"

OpenStudy (anonymous):

so not a carat